### v1

# Deep-Learning-Driven Prefetching for Far Memory

Abstract.
Modern software systems face increasing runtime performance demands, particularly in emerging architectures like far memory, where local-memory misses incur significant latency. While machine learning (ML) has proven effective in offline systems optimization, its application to high-frequency, runtime-level problems remains limited due to strict performance, generalization, and integration constraints. We present FarSight, a Linux-based far-memory system that leverages deep learning (DL) to efficiently perform accurate data prefetching. FarSight separates application semantics from runtime memory layout, allowing offline-trained DL models to predict access patterns using a compact vocabulary of ordinal possibilities, resolved at runtime through lightweight mapping structures. By combining asynchronous inference, lookahead prediction, and a cache-resident DL model, FarSight achieves high prediction accuracy with low runtime overhead. Our evaluation of FarSight on four data-intensive workloads shows that it outperforms the state-of-the-art far-memory system by up to 3.6 times. Overall, this work demonstrates the feasibility and advantages of applying modern ML techniques to complex, performance-critical software runtime problems.
1. Introduction
Machine learning (ML) has recently demonstrated strong potential in addressing a variety of systems challenges. Prominent examples include optimizing operating system and virtual machine configurations (yang2022stun, ; cheng2024dynamic, ; cano2018adares, ), predicting request arrival times for cluster-level resource management (resource-central, ; ml-datacenter, ; peng2019dl2, ; Paragon, ), and estimating object hotness to improve data placement in heterogeneous memory (ML-Object-HeteroMem, ; lu2024hammerefficienthotcolddata, ; IDT-RL-TieredMem, ). These successes typically rely on models that are either trained and evaluated offline or invoked infrequently at runtime—such as during object allocation. However, the performance of applications is directly influenced by frequent, fine-grained events like memory accesses and instruction execution.
So far, ML-based solutions for such high-frequency events are largely confined to the micro-architectural level. Examples include CPU branch prediction (TinyBERT-BP, ; tarsa2019-bp, ), instruction prefetching (pierce1994wrongpath, ; ramesh2024pdip, ; protean2023, ), and cache line prefetching (duong2024twilight, ; zhang2023dart, ; voyager, )). Unlike hardware-level ML techniques that operate on small, consistent inputs with tight latency and resource constraints, applying ML at the software systems runtime layer introduces distinct challenges and opportunities. At the systems level, application behavior can vary widely by input data, application configurations, and environments, and application performance is impacted by longer-term history. Meanwhile, system software has more flexible access to idle CPU cores and memory space, potentially allowing for more resource-demanding ML techniques. This discrepancy raises a fundamental research question:
How to properly leverage ML techniques to solve complex, performance-critical runtime problems at the systems layer?
To answer this question, we explore a representative and challenging problem: ML-based far memory prefetching in the Linux kernel. Far-memory systems—where most of the data resides in remote memory and only a small subset is cached locally—are gaining traction for reducing overall memory costs through the use of cheaper, network-attached memory and memory sharing across nodes. In far-memory systems, application performance is highly sensitive to the latency of fetching data on demand from remote memory, as a far-memory access over RDMA takes more than 20 times of a local DRAM access.
To make far memory viable, the on-demand data fetching overhead must be mitigated. Today’s predominant approach is to improve far-memory communication bandwidth and latency with interconnect technologies like CXL (cxl_spec2023, ; chen2023pond, ; wang2023dmemos, ; subramaniam2022cxl, ; chen2024nextgen, ). However, these approaches require substantial hardware upgrades and ecosystem adoption, slowing their deployment. An alternative and orthogonal strategy is to reduce the frequency of on-demand fetches through software prefetching. Today’s far-memory systems employ rule-based prefetchers (e.g., sequential or strided access patterns (mira, ; almaruf2020leap, )), which perform well for regular memory access patterns but fail in the face of complex access behaviors influenced by non-trivial interactions between application phases, control flow, data locality, and cache hierarchy.
Many data-center workloads including graph processing (PageRank, ; han2024graph, ), tree and index structures (guttman1984r, ; gusfield1997algorithms, ), pointer chasing (hsieh2017implementing, ), and recursive data structures (harold2004xml, ) exhibit memory access patterns that defy rule-based prefetching. If these access patterns could be learned and predicted accurately, far-memory systems could proactively fetch data and mitigate the performance penalties associated with remote access, even in the absence of new hardware.
Figure 1. FarSight Achieving Three Key Goals Together. FarSight, FastSwap (fastswap, ), and Hermit (hermit, ) are far-memory systems that run in the Linux kernel. Voyager (voyager, ), Hashemi etal. (pmlr-v80-hashemi18a, ), and Twilight (duong2024twilight, ) are micro-architecture CPU cache prefetchers implemented in simulation or with offline traces.
Modern deep-learning (DL) techniques such as LSTM (hochreiter1997long, ), RNN (socher2011parsing, ), and Transformer (vaswani2017attention, ) have demonstrated the ability to model complex, long-range, and nonlinear patterns that traditional heuristics cannot capture. However, incorporating such models in performance-critical system settings introduces a unique set of challenges. To ensure good application performance, ML techniques must impose minimal overhead. This makes online training impractical, as it introduces significant memory and computational costs. Offline training, while avoiding these runtime expenses, raises a different challenge: ensuring the trained model remains effective on new or unseen inputs. At runtime, predictions must be both accurate and lightweight–—incorrect predictions lead to prefetching irrelevant data, which wastes local memory and degrades application performance. Although larger models and longer input histories can enhance prediction accuracy, they also increase the cost of access tracking, inference latency, and memory usage due to the size of model weights. Ultimately, reconciling the competing goals of high accuracy, low overhead, and strong generalization remains a fundamental and difficult challenge (Figure 1).
To confront this challenge and demonstrate the benefits of ML techniques for performance-critical systems problems, we build FarSight, a Linux-based far-memory system that incorporates modern DL techniques for effective data prefetching. Our key insight is that memory access behavior is governed by both application semantics (e.g., algorithmic logic) and input-dependent runtime context (e.g., memory layout). The semantics tend to generalize across inputs and can be learned offline, but the actual memory addresses are input-specific and best handled at runtime. FarSight exploits this separation by training a DL model to learn semantic patterns and delegating address resolution to a runtime system component.
Specifically, we propose to represent application semantics as relationships between memory accesses. For each access, we observe that the subsequent access usually only has a small set of possibilities. We assign each possible outcome an ordinal identity. While the actual memory addresses corresponding to these possibilities vary across different inputs, the transition pattern—i.e., which ordinal is likely to follow given the history–—is often learnable and generalizable. For instance, in a linked list traversal, each accessed node is always followed by the next node in the list. Although the concrete addresses of these nodes differ per execution, the access behavior remains the same. Similarly, in the PageRank algorithm, a page’s rank update depends on the ranks of linking pages. The specific memory addresses of those linked pages depend on the input graph, but their identities can be precisely captured at runtime and predicted using ordinal labels.
To formalize the above idea, we set the DL model vocabulary as the anticpated outcome possibilities (i.e., a configurable $K$ defaulting to 64). The DL model uses memory access history sequences encoded into $K$ vocabulary and predicts future memory accesses as a sequence of 0 to $K-1$ ordinals. These ordinals are resolved at runtime via a lightweight future map–—a data structure that records actual addresses observed during execution. This design significantly reduces model vocabulary from the full memory address space to a small, fixed size, enabling high prediction accuracy with a compact DL model.
Building on top of this idea, we optimize FarSight’s performance while delivering high prediction accuracy with several techniques. (1) Asynchronous prediction and prefetching: Prediction is triggered when application threads block on far-memory fetches, utilizing otherwise idle CPU cycles and hiding inference latency behind the critical path. Afterward, prefetching requests are executed asynchronously in the background. (2) Multi-step lookahead: The model predicts several steps ahead, enabling timely prefetching before the application issues the corresponding memory accesses. (3) Compact DL architecture: We employ Retentive Network (retnet, ), a compact RNN-based model small enough to fit within a core’s L1 cache, ensuring low latency and energy efficiency. (4) Efficient input encoding: We use a position encoding scheme that supports reuse of cached context across predictions, reducing redundant computation and memory access overhead.
We implement FarSight’s training as an offline process when an application is deployed. We implement FarSight’s runtime system in the Linux kernel, with most of it being a Linux kernel module and the rest changing Linux’s swap system slightly. We evaluate FarSight with four real-world applications and benchmarks: MCF (spec2006_429mcf, ), PageRank and Shortest Path from the GAP benchmark suite (beamer2015gap, ), and XGBoost (chen2016xgboost, ). We compare FarSight to FastSwap (fastswap, ) and Hermit (hermit, ), two Linux swap-based far memory systems. Our results show that FarSight outperforms FastSwap by up to 3.6 times and Hermit by up to 2.6 times.
Overall, this paper makes the following key contributions.
  • The first ML-based far-memory system, FarSight.
  • Three insights of opportunities and challenges in ML-based prefetching.
  • The novel idea of decoupling memory-access semantics and memory address layouts.
  • Full implementation of a DL-based swap system in the Linux kernel.
  • Various optimizations at the systems and prediction-method layers.
We will open source FarSight upon the paper’s acceptance.
2. Motivation and Related Works
This section first provides an overview of far-memory systems and far-memory prefetching. We then discuss why and when ML techniques are a good fit for far-memory prefetching. Finally, we discuss the challenges of applying DL techniques for far-memory prefetching.
2.1. Far Memory and Existing Prefetch Approaches
Far-memory systems are systems where applications have access to memory beyond CPU-local memory, e.g., memory at another server or memory in a disaggregated memory pool. Far memory allows applications to access larger amounts of memory and more efficient cluster memory utilization. For applications to utilize far memory, there usually is an indirection layer (e.g., a swap system (fastswap, ; hermit, ) or a user-space library (aifm-osdi20, ; mira, )) that fetches data from far to local memory.
The main limitation of far-memory systems is the communication delay between local and far memory. For example, for a local memory size that is half of the far memory size, naive implementation of a far-memory system could result in half of the accesses going to far memory, resulting in an application slowdown of 13 times for RDMA-based far-memory systems. In practice, to strive for higher memory resource efficiency, local memory sizes are often set to below half of the far memory size (chen2023pond, ).
To hide this delay, most far-memory systems prefetch future accesses from far memory and cache them locally. Existing far-memory systems (fastswap, ; aifm-osdi20, ; almaruf2020leap, ) prefetch far-memory data with rule-based approaches by detecting and following linear and strided patterns. As such, they are limited to only benefit applications with such regular memory access patterns.
2.2. Why and When ML for Far-Memory Prefetching
A significant number of data-intensive applications have memory access patterns that simple rules cannot capture. For example, graph-based algorithms like PageRank (PageRank, ) and graph search have access patterns that are highly dependent on graph structures. As another example, ML algorithms have memory access patterns dependent on ML model architectures. Figure 2 visualizes memory accesses in PageRank using the Twitter graph with 41M nodes (snap-twitter2010, ) and insertion sort of 1 million randomly generated values in a linked list. As seen visually, these applications follow specific memory-access patterns, but they are hard to describe by simple rules. This is because these applications have repeatable program logic that is not completely random or completely rule-based.
More generally, we expect ML techniques to work for applications with code pieces that are repeatedly executed (e.g., a loop, a recursive function, etc.) or follow some traversal behavior (e.g., link-list walk, graph walk, indirect array accesses, array accesses according to some algorithm, etc.). Many data-center applications, such as data analytics, ML algorithms, graph processing, tree-based indexes, sorting, etc., fit these features. On the other hand, memory accesses that are completely random or follow simple rules do not fit ML-based approaches; for these applications, FarSight falls back to using rule-based prefetching or no prefetching and thus performs similarly to existing far-memory systems.
Figure 2. Example Memory Access Sequence over Time Captured by tracing page access sequence during execution.
Insight 1. Many data-center applications show repeatable, patterned memory-access behaviors that can potentially be captured by ML but not by simple rules.
2.3. Challenges of DL for Far Memory Prefetching
Successful application of DL for far memory prefetching presents several unique challenges.
Performance and why not GPU. Unlike many existing exploration of ML for systems, prefetches are on the performance-critical paths that directly impact application performance. Prefetched data that arrive later than the on-demand accesses make them useless. Pausing applications to perform prediction and prefetching is also not an option, especially when they take long.
To put things in perspective, we measure model prediction time (also known as model forwarding time) on GPUs. Predicting one token with 64 input tokens using the Llama-7B model on an Nvidia A100 GPU takes around 15 milliseconds; even small models like NanoGPT-10M take hundreds of microseconds. To understand where the latency comes from, we build a minimal “model” with only one matrix multiplication and memory copying from/to the CPU. With this, we find that the launching time for the matrix multiplication and memory copy kernels takes close to 10 microseconds. In comparison, a far-memory 4 KB page read takes around 2 microseconds with today’s InfiniBand-based network.
The long delay of GPU-side model forwarding implies that the model needs to predict far into the future. This is because the memory access history that the model uses is before the forwarding starts, and by the time forwarding finishes, hundreds of thousands of memory accesses could have happened. Effective prefetching requires a prediction of at least this many accesses in the future, but long predictions are usually less accurate.
Insight 2. GPU-based memory-access prediction is slow and thus requires predicting farther into the future, making accurate prediction hard.
Accuracy. Although wrong prefetches do not affect application execution correctness, they waste local memory space and network bandwidth, which are especially precious under far-memory environments. As local memory is expected to run to capacity, a prefetched page will need another local page to be swapped out on demand, taking about 4 microseconds from our evaluation. Thus, it is essential to design ML techniques for accuracy. Model accuracy depends not just on model size but also largely on the prediction problem formation. For example, longer input history usually helps model prediction to be more accurate, but at the expense of higher monitoring overhead. As another example, a small vocabulary (i.e., number of distinct prediction values) is easier for a model to predict accurately, but makes it challenging to represent a complex problem space like memory addressing.
Generalization. Because of the performance requirement, it is infeasible to train or fine tune a model during the run time. An offline trained model avoids any runtime performance overhead but has no access to runtime status. Although many data-center applications follow memory-access patterns, the exact pattern is often input-dependent. Moreover, the actual memory addresses being accessed can be different across runs even with the same input because of memory address randomization techniques like ASLR.
Insight 3. It is challenging to achieve performance, accuracy, and generalization together, as they have competing goals.
3. FarSight Design
FarSight is a swap-based far-memory system that prefetches memory pages from network-attached far memory with an DL-based memory access predictor. FarSight consists of an offline training component and an online predictor and prefetcher component sitting in the Linux kernel’s swap system, as illustrated in Figure 3. We choose to build FarSight as a Linux-kernel-based swap system, as it allows all Linux-compatible applications to transparently use far memory (fastswap, ; hermit, ; canvas, ; almaruf2020leap, ). The core of FarSight can also be implemented in the user space.
When an application is deployed, FarSight trains a small model (3K-parameter Retentive Network (retnet, ) architecture) by tracking the execution of the application with user-supplied sample application inputs. During the run time, FarSight loads the trained model onto each CPU core running the application. FarSight predicts future far-memory accesses using its captured recent program execution history and issues the corresponding prefetch requests. When the prefetch is correct, future access will be local. Otherwise, or when prefetch is slower than the future access, the application thread issues an on-demand far-memory access.
This section discusses the core ideas and technical details of FarSight.
3.1. Two Key Ideas
Based on our insights from §2, we propose two primary ideas for FarSight.
Figure 3. FarSight Overall Architecture All red parts are FarSight.
Pattern and addressing decoupling. The first core idea revolves around how we frame the prediction task. We observe that an application’s memory access patterns often exhibit repetition due to constructs like loops and recursion. However, these patterns are typically complex and nonlinear, making them difficult to model with simple rules or heuristics. DL models, on the other hand, are well-suited to capturing such long-range, nonlinear dependencies. Nevertheless, to maintain efficiency—particularly with smaller models that can run within a single CPU core and are trained offline—we must simplify the problem space and minimize runtime variability.
Our idea is to decouple application memory access semantics from the actual runtime memory layout by using DL prediction for the former and mapping tables for the latter. Specifically, we use a small DL model to predict memory access relationships in terms of abstract ordinals—–representing the possible memory-access outcomes after a short history of memory behavior—–rather than concrete memory addresses or offsets, which are highly input- and environment-dependent. At runtime, we construct future maps: mapping tables that resolve these predicted ordinals to actual memory addresses observed during the program’s first access, thus capturing the true memory layout dynamically.
In contrast to prior approaches that rely exclusively on ML prediction (§5) or purely on runtime history (as in traditional rule-based prefetching), FarSight integrates both DL prediction and runtime recording. By structurally decoupling the problem space, each component operates where it is most effective.
CPU prediction and I/O overlapping. Our second main idea focuses on the efficient realization of the DL-based prediction framework. As discussed in §2, GPU-based prediction has performance and accuracy issues. Thus, FarSight performs its model prediction in CPU. To avoid the energy and performance cost of additional CPU cores, FarSight performs prediction for an application on the same core it runs on. Our idea to avoid application performance overhead is to hide model prediction behind far-memory I/O time. Specifically, FarSight performs prediction when a foreground page fault is being handled with far-memory data read. We then issue asynchronous prefetch requests in the background.
Figure 4. FarSight Prediction Representation An example of vocabulary size ($K$) being 4. The top part shows code/algorithm corresponding to the accesses of chunks addr-x, addr-b, and addr-e. The bottom shows the input to the model: the chunk addresses and PCs of the 5 previous misses.
3.2. Prediction Task Formation
FarSight aims to reduce local memory misses by prefetching memory pages that are likely to be accessed in the near future. To reduce runtime overhead and improve the accuracy of a small DL model, we use a small vocabulary size of $K$, defaulted to 64. This means that both model input and output can only have $K$ values, implying that a future map should only have $K$ entries. Below, we explain how we achieve this small vocabulary size with our patten-addressing decoupling idea.
Model inputs. FarSight uses page miss history as the input to the DL model instead of full memory access history. This is because, by being in the swap system, FarSight can observe and log miss addresses on every page fault without incurring additional overhead. In contrast, capturing the full memory access stream would introduce substantial runtime overhead and is therefore avoided. In addition to using page miss addresses, we associate every miss with the faulting program counter (PC), as doing so can incorporate program execution information with memory access history, and recording and using PC incurs no additional overhead.
To fit the two types of inputs into the vocabulary, we take the mod of their value to the vocabulary size, $K$. We then use a history sequence of $h$ pairs of the modulo of miss page address and PCs as the model input, as shown at the bottom of Figure 4. Even though taking a mod is inevitably a lossy process, a history sequence and two types of information still allow our model to make accurate predictions.
Model outputs and future maps. We choose to predict page misses (i.e., accesses to memory pages not in local memory), rather than attempting to predict every individual memory access–—which would be computationally intensive and unnecessary. Essentially, FarSight uses page miss sequence in recent history to predict page miss sequence in the future. This approach significantly reduces the computational load on the DL model and the monitoring overhead.
A straightforward way to model memory miss prediction is by using their memory addresses, as used by most prior ML-based memory access prediction works (voyager, ; zhang2023dart, ; pmlr-v80-hashemi18a, ; peled2015semantic, ; peled2019neural, ). While straightforward, address-based prediction requires a huge vocabulary—$2^{3}6$ for 4KB pages in 64-bit systems. In comparison, English vocabulary used by modern Large-Language Models like GPT is only 50K to 100K in size (radford2019language, ; kaisugi2023gpt4vocab, ), beyond which prediction accuracy starts to degrade even for large models. Clearly, the huge memory address vocabulary does not meet far-memory prefetching’s accuracy demands (§2).
Our solution is to label possible outcomes of memory access as ordinals. Specifically, we record a vocabulary size (i.e., $K$) of possible next memory page misses after a miss happens at page $X$. Based on the model inputs as described above, our model predicts an ordinal from 0 to $K-1$, corresponding to one of the likely next page misses. We dynamically maintain a future map) for each page $X$. Each entry in the future map represents one possible page to be accessed after the miss of page $X$, and the value of the entry is the runtime virtual memory address of the page. A null future map entry represents an outcome that has not occured at the runtime yet. We delay the detailed discussion of setting up and maintaining future maps to §3.3.
Figure 4 illustrates this idea with an example. Each page is associated with its future map of size 4 (i.e., a vocabulary size of $K=4$). For the page of addr-x, its next access can be any of the four cases in the switch clause, and the model predicts one of them based on the access history. At the state of Figure 4, only addr-y have been accessed by the application; i.e., the model has previously predicted the ordinal $1$ a fault on addr-x happened, and the subsequent access of that addr-x fault is addr-y. Similarly, on the right-hand side, pages addr-b and addr-e are tree nodes and are accessed by a Breadth-First-Search program; their next accesses are their child nodes.
Vocabulary size. Naturally, a program can have fewer or more possible memory pages to access than $K$ after a page is accessed. If there are fewer possibilities (e.g., pages addr-b and addr-e only have three and two possible outcomes), the model will just not yield the remaining as a predicted value. If there are more possibilities than $K$, the model will not properly capture the less frequently occurring accesses. We will detail how we train the model so that it predicts the $K$ most frequent next accesses in §3.5.
Figure 5. FarSight Threading Model
We set the configurable $K$ to 64 by default, which strikes a balance of memory overhead and prediction accuracy. Note that $K$ outcomes represent $K$ 4 KB pages, which contain a $4K$ KB address range (256 KB by default). The default value of 64 works well for all our applications for a few reasons. First, small future maps allow for hot entries to be cached at CPU L1 and L2 caches, largely reducing the prediction latency. As will be discussed in §3.3, the default 64 value already necessitates the need for swapping cold future maps to far memory when local memory is small. Second, applications with repeatable behavior usually have limited possible outcomes after one page fault. For example, pointer chasing, database B-trees, common program control flows, and sorting algorithms have one to a handful of possible memory outcomes. On the other hand, a graph with high skew could have some nodes with a large number of neighbors, but the frequency of accessing these neighbors is relatively low, and failure of prefetching them does not largely impact application performance, as shown by our PageRank results (§4.1). Third, most allocators assign addresses from a range to closely requested address allocations, resulting in most accesses being within the same memory page and $K$ pages being able to host them.
DL model architecture. We adopt the Retentive Network (RetNet) model architecture (retnet, ), which unifies the benefits of Transformer (vaswani2017attention, ) and RNN (socher2011parsing, ) to achieve O(1) inference latency and O(N) inference memory space, where N is the sequence length, while maintaining good accuracy and training speed. It achieves this goal primarily by replacing the Transformer’s softmax operation with a weighted sum of the sequence’s history context, as shown in Figure 7. By applying an exponentially decaying factor to the context, it gives more weight to the recent history (i.e., attends more to recent tokens). Even though this generic decaying method may not work with all natural language cases, it fits our usages, as program behaviors are usually influenced more by recent history than distant history. More importantly, its superior inference latency and memory consumption allow us to deploy it on each CPU core (§3.3). We feed the memory address (after mod $K$) as Q and PC mod $K$ as K and V.
3.3. Thread Model and Metadata Management
Before delving into our optimizations to the prediction method 3.4, we first discuss how we efficiently integrate model prediction, prefetching, and future-map maintenance while ensuring minimal impact to foreground application performance.
CPU-based, I/O-overlapped model prediction. As discussed in §2.3, prediction on GPU is not a viable option for far-memory prefetching. Our approach is to utilize the CPU core of each application thread to perform prediction while the thread waits for far-memory I/O operations. We observe that even with effective prefetching, on-demand network accesses (misses that are not prefetched) still occur occasionally when the local memory size is small or when applications do not exhibit good locality. When an on-demand page miss occurs during a page fault, the application thread must wait for the missed page to be fetched from main memory before proceeding. We leverage this otherwise idle CPU time to perform model prediction. Doing so avoids the latency of switching to a different CPU core and eliminates any synchronization needs across different cores, making it more scalable.
Furthermore, we maintain all model weights, application memory access history (model inputs), and hot future map entries in the core’s L1 and L2 caches. This is feasible because of the small model size and small history window we choose (totalling 20 KB of weights and metadata). As such, one model prediction takes less than 600ns, significantly faster than a network round trip of around two microseconds with RDMA.
After the model prediction returns, FarSight issues an asynchronous prefetch I/O request to far memory and immediately yields the CPU core to the application thread. Prefetched pages are placed in the Linux swap cache.
Future map management. A future map is created and associated with a memory page when the page is first accessed (and thus causing a page fault). Initially, all elements in the future map are null, implying that the system does not yet know about memory addresses for future accesses. After the model makes a prediction at the time when memory page $X$ is faulted, we look up the future map of $X$ with the index of the predicted ordinal $k$. If the $k$th element in the future map is not null, FarSight issues an asynchronous prefetch request with a far-memory address as the $k$th element’s value. If the $k$th element is null, we do not perform any prefetching, as the far-memory address has yet to be determined (this is the first-time access). When the prefetched page is being accessed, it means a page fault happens with a hit in the swap cache, and page is moved from the swap cache to for regular memory accesses. At this faulting time, we update the $k$th element’s value in page $i$’s future map with the faulting page virtual memory address.
Figure 6. Indirected and Swappable Future Maps. Application pages have one-to-one mapping to future-map roots based on their virtual page number (VPN), while the mapping from future-map roots to future maps is dynamic. On the left, the blue page (VPN 3) and its associated futuremap are swapped out. On the right, the yellow page (VPN k) and its futuremap is swapped in.
With a default $K$ value of 64 and each future map entry taking 4 bytes, each 4 KB page, including those residing in far memory, requires 256 bytes, or 6.25% of memory space. When the local memory size is small, future maps can consume significant space if all are stored in the local memory. For instance, with a local memory of 20% total memory size, future maps consume 6.25% of total memory, or 31.25% of the local memory. If this significant portion of local memory cannot be used to store application in-memory data, application performance will be largely impacted.
We propose two solutions to resolve this problem, as illustrated in Figure 6. First, we make future maps swappable at the granularity of one single future map (256B for $k=64$). When an application memory page is swapped out, we also swap out its associated future map to far memory. When an application page is swapped in at prefetching or on-demand faulting time, we also swap in its future map. Second, we add a level of indirection to support fine-grained future-map memory management—a small number of non-swappable future-map root pages. Every application page has a static mapping to a future-map root based on its virtual page number. Each future-map root stores the current virtual address of the application page’s associated future map. As future maps are swapped out, new ones take their locations by establishing new maping between roots and future maps. Doing so avoids internal fragmentation and improves the memory efficiency of future maps stored at the local memory.
3.4. Prediction Method Optimization
The previous two sections details our prediction problem formation and system implementation methods. We now describe two additional optimization methods we take to further improve FarSight’s overall performance, as shown in Figure 7.
Figure 7. FarSight’s Prediction Optimization Methods Demonstrating the use of each history window to predict $s$ misses ahead of time and predicting $f=2$ pages at a time.
Multi-step lookahead. So far, we assume the model predicts the immediate next missed page. With such an approach, even if we issue a far-memory access request right after the prediction, the communication delay is likely longer than when the next miss happens, making the prefetched data arrive too late to be useful. Our solution is to predict farther into the future with a look-ahead distance to cover the communication delay to prefetch application data. Specifically, we predict and prefetch the $s$th future memory miss from the current access (i.e., $s$ is the look-ahead distance). We determine $s$ by conservatively choosing a large percentile (e.g., 95%) of profiled communication delay distribution, $d$, and the average profiled inter-arrival time between two memory accesses, $l$; $s$ is $d/l$. With this conservative setup, prefetched data could arrive before it is needed but rarely after. Furthermore, to efficiently utilizing network bandwidth, we prefetch $f$ pages at a time. To know what $f$ pages to prefetch, we leverage the autoregressive nature of the RetNet (retnet, ) model to predict $f$ future misses in a sequence at a time.
Model input encoding. At the time of a miss, we use the recent history window of $h$ misses as the model input sequence. A naive way to encode the history is to treat each miss as one token and perform positional encoding of these tokens starting from position 0. This encoding works for generic sequence-to-sequence problems, as each new request to the model is treated as a different sequence. However, in our environment, most tokens (past accesses) but one overlap when we move the history window by one position ($m_{3}$ to $m_{h}$ accesses overlap between the two predictions shown in Figure 7). With the naive encoding method, we could not reuse any previously computed intermediate results as the token positions have changed in the new window, which causes recomputation performance overhead.
To solve this problem and improve prediction performance, we propose a new encoding method based on rotary positional encoding (su2021roformer, ), a widely used encoding algorithm. We observe that the positional relationship between tokens is important but not the absolute starting position. Instead of always starting from 0, our encoding starts from the position where the first access in the current history window (e.g., $m_{3}$ in t2 window) was at in the previous window ($m_{3}$ at the 60-degree angle). Essentially, we turn the rotary wheel by one unit of angle at every prediction step to align the same access at the same angles. This allows us to reuse the computed context of overlapped accesses (e.g., $m_{3}$ to $m_{h}$).
3.5. Model Training
Each deployed application goes through an offline training process. We start the process by executing the application with a user-supplied sample input with fully local memory on a single server. We expect the sample input to be smaller than the actual runtime inputs and can thus run fully locally. We do not run the training sample run in a far memory setup as we expect one trained model to work across different far memory settings (e.g., different local memory sizes, different network speeds). As will be shown in §4.1, a model trained with small inputs generalizes well to different larger inputs thanks to FarSight’s decoupled representation. To adapt to different inference far-memory settings, we randomly drop out (exclude) a certain percentage (e.g., 10%) of memory accesses.
We instrument the sample run to capture all memory accesses and then train the RetNet model with this collected trace in an offline manner. The training process uses the same vocabulary size, look-ahead distance, encoding method, and history length as introduced in §3.2 and §3.4. As the sample run executes fully locally, there is no miss or prefetching. Thus, we build a future map for each accessed memory page to track its $s$th subsequent access. As we have the oracle knowledge of the whole execution, we maintain the top $K$ most frequently accessed subsequent pages in each future map. The training target is the correct index in the future map that matches the ground truth (with full trace, we have the oracle of what page will be accessed next).
4. Evaluation Results
Figure 8. MCF Latency. Higher is better.
Figure 9. XGBoost Latency. Higher is better.
Figure 10. GAP PageRank Latency. Higher is better.
Figure 11. GAP SSSP Latency. Higher is better.
Figure 12. MCF Prefetch Hits. Higher is better.
Figure 13. XGBoost Prefetch Hits. Higher is better.
Figure 14. MCF Prefetch Hit Rate.
Figure 15. XGBoost Prefetch Hit Rate.
This section presents our evaluation results of FarSight.
Implementation. We implemented FarSight with 5.5K lines of source code in the Linux kernel. FarSight currently runs on one compute and one memory node configuration, but can apply to memory pooling with multiple memory nodes. For consistent and fast inference, we implemented the inference of RetNet with AVX512, a CPU vector instruction set available in the x86 platform. Current implementation supports a max sequence length of 64 but is adaptable to higher or lower sequence lengths given different computation latency requirements. The detailed hyper-parameters of models can be found in Table 1.
Environments. We evaluate FarSight on our private clusters. The compute node is running on a server equipped with a 28-core Intel Xeon Gold 5512U CPU (2.1 GHz) and 16 GB RAM. The memory node is running on a server equipped with a 16-core Intel Xeon Gold 5218 CPU (2.3 GHz) with 64 GB RAM. Both servers are connected with 100 Gbps Mellanox EDR-CX4 NIC through a 100Gbps RoCE ToR switch.
Parameter Value
Number of parameters 2240
Hidden dimension ($d_{\text{model}}$) 8
Number of attention heads ($d_{\text{attn}}$) 4
Number of layers ($n_{\text{layer}}$) 2
Maximum sequence length ($T$) 64
Batch size ($B$) 1024
Learning rate ($\eta$) 0.003239
Loss function Cross Entropy
Optimizer AdamW
Table 1. Model configuration and training hyper-parameters.
Baselines. We compare FarSight with two baselines: FastSwap (fastswap, ) and Hermit (hermit, ). FastSwap is a swap-based far-memory system implemented in the Linux kernel. Hermit builds on top of FastSwap and improves its swap-out procedure to avoid swap-out being the application performance bottleneck. Both systems use the Linux prefetching policy, which only follows simple and strict rules for issuing prefetches for sequential accesses. We do not have any ML-based prefetching baseline, because existing ML-based prefetching solutions are all at the micro-architecture level and do not work on a real software systems like Linux (§5).
Workloads. We evaluate FarSight and baselines with four workloads, XGBoost (chen2016xgboost, ), PageRank and shortest path in GAP benchmark suite (beamer2017gapbenchmarksuite, ), and MCF (spec2006_429mcf, ).
XGBoost is a machine learning framework for gradient boosted decision trees (GBDTs). During training, it builds an ensemble of decision trees, where each successive tree attempts to correct the errors based on a specified loss function. Because of the irregular and data-dependent nature of tree traversal, it benefits from an ML-based prefetching. We run the HIGGS dataset (higgs, ) on XGBoost for a binary classification task that consumes 7.4 GB of memory.
The GAP Benchmark Suite is a collection of graph processing benchmarks designed to evaluate the performance of graph analytics systems. In our evaluation, we select two widely-used graph algorithms, Shortest Path (SSSP) and PageRank, which are commonly used in large-scale systems. These algorithms also exhibit complex memory access patterns that are challenging for rule-based prefetchers to capture. We use the suite’s built-in graph generator to create graphs ranging from 4 million to 16 million nodes. The memory consumption of these two workloads ranges from 2.2 GB to 8.2 GB.
MCF (Minimum Cost Flow) is one of the workloads from the SPEC-CPU2006 benchmark (spec2006_429mcf, ). Due to the intricate interconnections between nodes and edges in a cost-weighted graph, MCF exhibits unpredictable and highly variable memory access patterns that challenge conventional prefetching strategies. This benchmark effectively tests our predictor’s ability to make accurate decisions under diverse and dynamic input conditions. In our evaluation, we present results using a configuration with 220 MB to 390 MB of memory consumption.
4.1. End-to-End Application Performance
We first present the end-to-end application performance and prefetching effectiveness with FarSight and the two baselines.
Application performance. Figures 11, 11, 11, and 11 present the end-to-end application performance of MCF, XGBoost, PageRank, and SSSP, respectively. For each set of experiments, we change the application server’s local memory size from 30% to 90% of the total application memory size (X axis) and measure the total application execution time (Y axis). For each result, we normalize the application execution time against that of running at full local-memory capacity, and higher Y-axis values are better.
Across all four workloads and input settings, FarSight consistently outperforms FastSwap and Hermit by up to 3.6 times and 2.6 times, respectively. Comparing across local-memory size settings, FarSight achieves greater improvements when the local memory size is small—an environment especially challenging but useful for far-memory systems. A smaller local memory size increases the likelihood of missed accesses, amplifying the performance impact of an effective prefetching policy.
Figure 16. Handling Input Variance in MCF. Three inputs used for MCF (220 MB, 240 MB, and 390 MB) tested on the same trained model.
Figure 17. Input Variation in PageRank. Three graphs used for PageRank (2.2 GB, 3.9 GB, and 8.2 GB) tested on the same trained model.
Figure 18. Performance Breakdown of FarSight using MCF with $30\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
FarSight outperforms the rule-based far-memory baseline systems because it is able to make more future access predictions and thus issue more prefetches. Moreover, the prefetches have high hit rates (i.e., highly accurate). We will present further analysis of these effects with additional evaluations shortly.
Among the four workloads, FarSight performs best on XGBoost, where our DL-based prediction framework effectively captures the tree traversal pattern; however, the heuristic-based strategies used in the baseline systems fail to do so. MCF is the hardest to prefetch among the four workloads because of its irregular graph-processing patterns, as can be seen from the overall lower performance when local memory is small. Nonetheless, FarSight manages to improve its performance by up to 2.6 and 3.6 times over Hermit and FastSwap. FarSight’s improvement is relatively lower with the two GAP workloads. Both GAP PageRank and SSSP represent the input graphs in matrices form and perform traversals on them: PageRank iteratively accesses the graph across multiple rounds, while SSSP performs a one-time traversal. Because of the compact matrix representation of graph, these workloads exhibit more sequential patterns that can be captured by the baselines.
Figure 19. Performance Breakdown of FarSight using MCF with $50\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
Figure 20. Effect of Vocabulary Size on End to End Performance.
Figure 21. Effect of Vocabulary Size on Prefetch Hit Rate.
Prefetch hits and rates. To understand the effectiveness of FarSight’s DL-based prefetching, we analyze the total number of prefetches that are accessed by applications (i.e., prefetch hit count) and the ratio of them to the total amount of prefetches (i.e., prefetch hit rate). Figure 15 and Figure 15 present the number of prefetch hits with different local memory ratios using the MCF and XGBoost workloads, respectively. Figure 15 and Figure 15 present the corresponding prefetch rates. We do not include Hermit in these figures, as their prefetch events are stopped when being determined as unuseful, thus having even fewer prefetch hits than FastSwap.
For MCF, FarSight issues 15.3×15.3$\times$15.3 × more prefetch requests and achieves 13.2×13.2$\times$13.2 × more prefetch hits on average compared to FastSwap. This shows that FarSight can uncover more application memory-access patterns and turn them into fruitful prefetches. Similarly, FarSight issues significantly more prefetch requests than FastSwap for XGBoost.
Interestingly, FarSight’s prefetch hit rate is lower than FastSwap’s, although still high, ranging from 67% to 97% for the two workloads. FastSwap’s high prefetch hit rate is because of its conservative prefetching policy—only when a sequential access pattern is detected will it perform prefetching. As sequential patterns are regular, FastSwap’s hit rate is consistently high. However, this conservative behavior results in FastSwap’s overall lower prefetch hit count and worse end-to-end application performance.
4.2. Deep Dive
We now perform a deep dive to understand and evaluate FarSight’s performance benefits.
Generalization to different inputs. To demonstrate our model’s ability to capture program semantics across program inputs, we evaluate the end-to-end application performance using multiple inputs with different sizes with the same model trained by another input. Figure 18 and Figure 18 show the results of these experiments. To challenge FarSight  we train the model on a small dataset and with larger, unseen inputs.
For MCF, we trained the DL model using a graph with 5K nodes and 50K edges. We test it with three different graphs, each with a distinct structure that is different from the training graph and contains between 20K and 40K nodes and over 200K edges. For PageRank, the model is trained using a graph with 10k nodes and 100k edges, and is tested on graphs with 4M, 8M, and 16M nodes, respectively. Both applications were tested under $30\%$ and $50\%$ local memory ratios.
FarSight demonstrates strong generalization capability with one-time offline training, effectively adapting to a wide range of input structures and runtime environments without the need for retraining or fine-tuning. In the case of MCF, this result in an average performance improvement of 2.3×2.3$\times$2.3 × over FastSwap, while in PageRank, we observe a 1.9×1.9$\times$1.9 × gain. FarSight is able to generalize across inputs and adapt to larger inputs with models trained on smaller ones, thanks to our memory-access behavior and address layout decoupling. These results demonstrate that effective models can be trained using short sequences of memory accesses as training data points, provided they reflect the application’s core memory-access patterns. This approach can substantially reduce both the training data collection overhead and the computational cost of model training.
Ablation study. To understand the performance benefits of each individual technique used in FarSight, we evaluate the system incrementally by adding one technique at a time, as shown in Figure 18 and Figure 21 using the MCF workload with 30% and 50% local memory. We present results under two different memory ratios to highlight how each technique performs under varying levels of memory pressure.
The leftmost bar in each figure represents a baseline configuration—the vanilla Linux setup without any prefetching. We then add FarSight’s DL-based prediction and trigger the prediction and prefetching synchronously on every page fault, including both faults that result in a local-memory miss and faults that hit the swap cache. Performing this basic DL-based prefecthing has a small improvement over the baseline.
We then change the prediction and prefetching to only be triggered on page misses, as shown by the third bars. Although this gives FarSight less chance to perform prediction, we can significantly avoid the runtime performance overhead that would otherwise be incurred for prediction on page hits. This is because a page fault that hits the swap cache will issue no far-memory I/O, leaving insufficient time for FarSight to finish its prediction.
So far, the prefetching FarSight performs is synchronous—application threads wait for it to finish. We then change the system to perform prefetching asynchronously, as shown in the fourth bars. FarSight optimizes latency by allowing the system to return to userspace as soon as the on-demand page is ready, rather than waiting for all prefetched pages to complete.
Next, we add the technique of swappable and indirected future maps. By proactively evicting cold future maps, FarSight frees up memory for user applications, leading to further performance improvements.
The final technique incorporated into the full FarSight design focuses on improving the timeliness of prefetching to avoid late prefetch—–a scenario where a page fault still incurs I/O latency even though the page was prefetched. By predicting and issuing prefetches earlier, FarSight ensures that prefetched pages are more likely to arrive before they are actually needed. As the latency results indicate, this strategy effectively reduces blocking and provides the final boost in system performance.
Sensitivity Tests By default, FarSight uses a vocabulary and future-map size of 64 to strike a balance between memory overhead and prediction accuracy. To evaluate how FarSight performs under different vocabulary size, we change $K$ from 8 to 64. Figure 21 and Figure 21. Overall, FarSight’s benefits prevail across $K$ values compared to FastSwap and Hermit, showing its robustness to sensitivity. FarSight’s performance degrades slightly when $K$ is smaller than 32. This is because a small future map size cannot capture the possible outcomes of application memory accesses. For example, consider an application with a recurring access pattern involving 1q frequently accessed pointers in an alternating sequence. If the future map size is limited to just 8 entries, the prediction pointers will overwrite each other, leading to reduced prefetch accuracy and a lower hit rate. This is evidenced by the decreased prefetch hit rates in Figure 21. Meanwhile, future map sizes larger than 64 add runtime performance overhead both becaseue future maps need to be swapped more frequently and because they can be less cached in CPU cache. Thus, we set the vocabulary size to be 64 by default, which works well for many typical applications.
5. Related Work
This section discusses related works in ML-based CPU cache and far-memory prefetching.
5.1. ML-Based CPU Cache Prefetchers
Prior research works have explored using ML techniques for local server prefetch predictions at the micro-architecture level by prefetching data from memory into CPU cache. However, because of their performance and/or accuracy issues, they have only been realized in simulation or for offline trace analysis. For example, Peled et al. have proposed reinforcement-learning-based (peled2015semantic, ) and regression-based (peled2019neural, ) approaches for memory prefetching. The former sets up the prefetching prediction as a classification problem and can only accommodate four possible address offsets for each prediction. The latter has accuracy issues as a regression model aims to be close to the ground truth, but correct prefetch requires the exact truth.
Another series of research works use LSTM to predict memory access sequences in the form of memory addresses (voyager, ; pmlr-v80-hashemi18a, ; srivastava2019predicting, ). The major problem with these solutions is their large vocabulary size, which results in slower prediction and low model accuracy (especially small ones). For example, no sequence-to-sequence models can handle a vocabulary size of $2^{48}$ — the possible addresses in a 64-bit address space for the current Intel CPU. To reduce the vocabulary size, Hashemi et al. (pmlr-v80-hashemi18a, ) limit the prediction to address deltas within a spatial region. However, this approach prevents predictions of data accesses that are distant from each other. Voyager (voyager, ) decouples memory pages and offsets within a page and only covers addresses that an application uses. A major issue with this approach is that a trained model with one application execution cannot be used by another execution, as different inputs or address randomization can all result in different sets of used addresses. As a result, Voyager requires an online model training every million memory accesses and “does not yet make neural models practical” (voyager, ).
More recently, Neural Network and Transformer-based models have also been applied to CPU cache prefetching (duong2024twilight, ; zhang2023dart, ). Twilight and T-LITE (duong2024twilight, ) use the combination of a customized two-layer neural-network model, clustering, and frequency-based history table for CPU cache prefetching. DART (zhang2023dart, ) distills a transformer model and then transforms the distilled model into a hierarchy of table lookups to reduce runtime performance overhead. Although these works have shown their CPU cache prefetch effectiveness through simulation, their training and prediction processes are complex and lack generalization or consistent accuracy.
Overall, unlike FarSight, these micro-architecture level CPU cache prefetching systems are unfit for real deployment in far-memory runtime systems, as they cannot achieve high prefetch accuracy, low overhead, and application generalization simultaneously. FarSight demonstrates the unique challenges and opportunities of applying ML techniques to software system problems, examplified by the challenging far-memory prefetching problem.
5.2. Far-Memory Prefetching
Existing far memory systems have explored various approaches for improving swap and data prefetching performance. On the mechanism side, Fastswap (fastswap, ) optimizes the swap datapath, and Hermit (hermit, ) further proposes asynchronous, proactive swap-out techniques. These systems improve execution efficiency, but still rely on the default Linux prefetching policy, which only follows simple rules and often fails to trigger under many real-world applications. Leap (almaruf2020leap, ) enhances the Linux prefetcher by attempting to match application page access patterns to a set of predefined trends, such as sequential and strided patterns. However, it struggles with complex, runtime-dependent application memory access patterns.
Different from swap-based far-memory systems, Mira (mira, ) is a fine-grained far-memory system that incorporates static program analysis, compiler optimization, and runtime profiling. Mira captures static program semantics and inserts prefetching operations with its compiler transformation. However, its performance degrades for applications that depend heavily on runtime information, such as graph-based workloads. Moreover, for applications to use Mira and other library-level far-memory solutions (aifm-osdi20, ), applications need to be ported with non-trivial manual effort.
Unlike FarSight, these prior systems fall short in far-memory prefetching scenarios, where access patterns are constructed dynamically at runtime, making them difficult to capture with pre-defined heuristics and highly sensitive to variations across application inputs.
6. Conclusion
We presented FarSight, a DL-based far-memory prefetching system in the Linux kernel. FarSight’s core idea is to decouple the learning of application semantics from the runtime capturing of memory accesses. By doing so and with our set of optimization techniques, FarSight achieves overall application performance benefits over two recent far-memory systems, by up to 3.6 times and 2.6 times. FarSight demonstrates the feasibility of deploying modern ML techniques to solve performance-critical problems in complex runtime systems. Future systems researchers and practitioners could leverage lessons we learned and building blocks of FarSight’s DL model, prediction problem presentation, and system-integration mechanisms.
Acknowledgements.
We would like to thank Ryan Lee, Geoff Voelker, Zijian He, Vikranth Srivatsa, and Reyna Abhyankar for their valuable contributions and feedback on this paper. This material is based upon work supported by funding from PRISM center (part of SRC’s JUMP 2.0) and gifts from AWS, Google, and Meta. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of these institutions.